# A Scoping Review of Synthetic Data Generation for Biomedical Research and Applications

### v1

@@ Results
Table 1: A summary of data types, generation Methods, and medical applications for our collected studies. We use language codes for each study’s data language, including English (EN), Dutch (NL), French (FR), Chinese (ZH), and Arabic (AR). The K (1,000) shortens spaces of the size column. We use acronyms for medical applications, including QA (question answering), IE (information extraction), and SDoH (social determinants of health). We use the “-” to indicate not specified information.
Study Modality Lang. Architecture Approach Size (K) Year Accessible Purpose Medical Application
[17] Text EN Transformer Fine-tuning 500 2020 No Training Biomedical QA
[18] Text EN Transformer Specialized model 11 2020 Yes Training Phenotype inference
[19] Text EN GPT Fine-tuning <1K 2021 Yes Training Clinical IE
[20] Text NL GPT, RNN Fine-tuning 355 2021 Yes Training De-identification
[21] Text EN GPT Prompting 45 2021 No Training Dialogue summarization
[22] Text EN GPT Fine-tuning 48 2022 No Training Readmission prediction
[23] Text EN GPT Fine-tuning 200 2023 Yes Training General
[24] Text FR GPT Fine-tuning 5 2023 Yes Training Clinical IE
[25] Text EN GPT Fine-tuning 5 2023 No Training Syndromic detection
[26] Text EN GPT Fine-tuning <1K 2023 Yes Training Assisted suicide
[27] Tabular EN Transformer Specialized model 976 2023 Yes Training General
[12] Text EN GPT Prompting 32 2023 No Training Alzheimer’s disease
[28] Text EN GPT Prompting <1K 2024 Yes Training Discharge summary
[29] Text EN GPT Prompting 2 2024 Yes Training PTSD diagnosis
[30] Image ZH, EN GPT Prompting 4,000 2024 No Training Medical imaging
[31] Text EN GPT Prompting 9 2024 Yes Training Acute renal failure
[10] Text EN GPT Prompting 4 2024 No Privacy Phenotype inference
[32] Text EN GPT Prompting 2 2024 No Training Cohort selection
[33] Text EN GPT Prompting <1K 2024 No Supplement Cohort selection
[34] Text DE GPT Prompting 4 2024 No Training Emergency medicine
[35] Text EN GPT Prompting 19 2024 Yes Training Psychological therapy
[36] Text EN GPT Fine-tuning 30 2024 Yes Training Medical QA
[37] Text EN Transformer Fine-tuning 2 2024 Yes Training Medical QA
[11] Text EN GPT Prompting 5 2024 Yes Training General
[38] Text EN GPT Prompting 2 2024 Yes Supplement SDoH extraction
[39] Tabular EN GPT Prompting <1K 2024 No Supplement Medical education
[13] Text EN GPT Prompting 30 2024 Yes Supplement Dialogue generation
[40] Text EN GPT Prompting 14 2024 No Training SDoH extraction
[41] Tabular EN GPT Prompting - 2024 No Training Gait rehabilitation
[42] Text EN GPT Prompting 158 2024 Yes Training General
[43] Audio EN GPT Prompting 87 2024 Yes Supplement EMS assistance
[44] Text EN GPT Prompting 37 2024 Yes Training Clinical IE
[8] Text EN GPT, T5 Fine-tuning 148 2024 No Supplement Suicide prediction
[45] Text EN GPT Prompting <1K 2024 Yes Training Cancer symptom extraction
[46] Text EN GPT Prompting 5 2024 Yes Training Clinical summary
[47] Text EN GPT Prompting 5 2024 Yes Training Medical QA
[48] Tabular - GPT Prompting <1K 2024 No Supplement Nursing care
[49] Text EN GPT Prompting <1K 2024 Yes Training Patient portal triage
[50] Text EN GPT Prompting 100 2024 No Training Report generation
[51] Text AR GPT Prompting - 2024 Yes Supplement Health chatbot
[52] Tabular EN GPT Prompting <1K 2024 No Supplement Clinical trial simulation
[9] Text EN GPT Prompting 3 2024 No Supplement Depression symptom
[53] Text EN Transformer Prompting 4 2024 No Supplement Autism diagnosis
[54] Text EN GPT Prompting <1K 2025 No Privacy Stroke prediction
[55] Audio EN GPT Prompting 3 2025 No Training Dialogue analysis
[56] Text EN GPT Prompting 459 2025 No Supplement Phenotype ontology
[57] Text EN GPT Prompting 5 2025 No Training Radiology prediction
[58] Text ZH GPT Prompting 270 2025 No Training Anesthesia QA
[59] Tabular - Transformer Specialized model - 2025 Yes Supplement Patient outcome prediction
[60] Text ZH GPT Prompting 6 2025 Yes Training Mental health IE
[61] Tabular EN GPT Prompting 6 2025 Yes Privacy Data scarcity
[62] Text EN GPT Prompting 88 2025 No Privacy Phenotype inference
[63] Text EN GPT Prompting 5 2025 Yes Training Medical reasoning
[64] Text ZH GPT Prompting 45 2025 No Training Database search
[65] Text EN GPT Prompting 74 2025 No Training Metastases detection
[66] Image EN Transformer Fine-tuning 2 2025 Yes Supplement Colonoscopy analysis
[67] Image EN GPT Prompting - 2025 No Supplement Pathology analysis
[68] Text ET GPT Fine-tuning 4 2025 No Training Clinical IE
[69] Tabular EN GPT Prompting 10 2025 No Training Patient outcome prediction
Synthetic Data Types and Medical Applications
Increasing model developments for biomedical applications lead to the surging needs of synthetic data [18, 20]. In this section, we provide an overview of recent studies using synthetic data and address three critical questions: 1) What types of synthetic data are being generated? 2) What modalities and clinical tasks do these datasets target? 3) Are those data resources accessible for reusable research, such as benchmarking and model training? To answer those questions, we examine several aspects of the collected studies and summarize them in Table 1, including data modality, data language, data size, published year, medical application, and data or model accessibility. We aim to provide insights into current and future trends of data types and biomedical applications.
Biomedical synthetic data has shown diverse medical topics, languages, and modalities. Most of the studies (55.9%, $n=33$) are for general topics, while others include specific medical areas, such as suicide [26, 8], colorectal cancer [66], and radiology [57]. For example, studies generate synthetic data for augmenting diagnosis accuracy for Post-traumatic stress disorder (PTSD) [29] and Autism Spectrum Disorders (ASD) [53]. English is the predominant language (84.7%, $n=50$), though there are also studies in French [24], Chinese [30, 58, 60, 64], German [34] and Arabic [51], reflecting some progress toward non-English-speaking patients. the study [34] deployed multilingual LLMs to generate German data to simulate realistic emergency-medical dialogues between ambulance staff and patients. Most datasets focus on unstructured text modality (e.g., clinical narratives [13, 21, 50] and discharge summaries [28]), while a smaller but growing subset includes tabular [27, 39, 41, 48, 52, 59, 61, 69], image [30, 66, 67], and audio [55]. For example, Theodorou et al. [27] generate novel tabular datasets of high-dimensional longitudinal EHR records to provide realistic, privacy-preserving alternatives for machine learning and statistical analysis, while Ejiga et al. [66] generate synthetic colonoscopy image datasets via fine-tuned text-to-image synthesis to augment training data for robust colorectal cancer detection and precise polyp segmentation. Data sizes vary across health issues (e.g., cancer and mental health), yet the majority has relatively small sets with fewer than 10K samples, indicating the critical utility of synthetic data under low-resourced scenarios. For example, FairPlay [59] generates synthetic data from authentic EHR records to bolster under-represented patient subgroups, boosting mortality-prediction F1 scores by up to 21% and markedly narrowing performance gaps across different demographic groups. The various synthetic datasets highlight a growing community to deploy the biomedical synthetic data.
Synthetic data generation in biomedical research increasingly supports a wide array of clinical applications. Those tasks cover diverse medical needs, including phenotype classification [45, 53, 25, 12, 65], PSTD symptom extraction [29], de-identification (SynNote) [20], clinical note summarization  [28, 46], and mental health assessment  [9, 35, 8]. For example, CALLM [29] constructs clinical interview datasets to aid PTSD diagnosis by generating synthetic transcripts; and Barabadi et al. [65] generated synthetic radiology reports to enable automatic detection of metastatic sites in cancer patients. Those applications broadening the utility of synthetic data for health challenges, such as data shortage [29] and privacy concerns [20, 10, 62, 61, 54]. However, accessibility remains a central concern for advancing synthetic data utility. Table 1 shows 50.8% ($n=30$) of the studies are explicitly described as accessible, providing open or partially open resources for reproducible development, while others’ accessibility details or licensing remain unclear. Open datasets such as Syn-HPI [19], ClinGen [11], and Syn-EMS-Audio [43] serve as synthetic data benchmarks for evaluating new models and methods, while others—particularly those derived from sensitive clinical domains—may be restricted or anonymized to protect patient privacy. Ensuring clear, well-documented access protocols and licenses will be essential for fostering broader collaboration and translational impacts of synthetic data in biomedical research.
Synthetic Data Generation Methodology
Emerging Transformer-based large language models (LLMs) have significantly advanced synthetic data generation methodologies in biomedical research, enabling diverse applications, as summarized in Table 1. These novel approaches address challenges associated with limited clinical data by augmenting datasets, thus enhancing generalizability and mitigating overfitting in downstream machine learning tasks across various healthcare domains. In this section, we systematically examine generation methodologies among 59 reviewed studies, summarizing recent trends in model architectures, generation techniques, and integration strategies for synthetic data generation. Specifically, we aim to answer two key questions: 1) what predominant methods (e.g., prompt-based vs. specialized architectures) are currently employed for biomedical synthetic data generation; and 2) what trends and variations do exist regarding model selection (open-source vs. closed-source) and prompting strategies (zero-shot, few-shot, knowledge-infused)?
Our review identifies prompt-based generation as the predominant synthetic data generation method, employed by approximately 72.9% ($n=43$) of the analyzed studies. Prompt-based approaches [64, 31, 52] primarily rely on meticulously crafted textual instructions, leveraging fine-tuned LLMs, such as T5 and GPT variants (e.g., GPT-4) for diverse biomedical tasks. Few-shot prompting provides a small number of curated examples (typically 2-5) within the prompt to illustrate the task and guide LLMs for synthetic data generation, which count 20.3% ($n=12$). The approach uses limited curated examples to guide LLMs in generating clinically relevant synthetic data across diverse applications, including suicide symptom extractions [8], phenotype classification [45, 53, 25, 12, 65], and therapeutic dialogue generation [13, 35]. For example, Ghanadian et al. [8] use a patient’s suicidal narrative and a medication counseling dialogue as few-shot prompts to generate clinical texts for suicide symptom extraction and therapeutic dialogue generation, while Nievas et al. [32] use clinical trial inclusion and exclusion criteria paired with patient summaries as few-shot prompts to generate a synthetic patient–trial matching dataset. Zero-shot prompting involves prompting LLMs without any illustrative examples and replies solely on explicit task instructions and the pretrained knowledge encoded in LLMs [61]. This approach is suitable for simpler tasks, such as dialogue summarization [21] or generating structured medical reports [50], where extensive contextual training is less critical. For example, Barr et al. [61] used zero-shot prompting with GPT-4o to synthesize perioperative clinical tables of patient demographics, surgical parameters and outcomes, showing that the model can produce realistic synthetic data and preserve key statistical patterns based solely on qualitative instructions. Additionally, ClinGen [11] studies utilize knowledge-infused prompting, which explicitly incorporates external biomedical resources, such as clinical ontology or structured knowledge graphs. The prompts integrate the biomedical resources to improve factual accuracy, clinical relevance, and contextual richness. For example, Chuang et al. [62] construct keyword-driven prompts by appropriate ICD-9 (International Classification of Diseases, 9th Revision) codes per primary diagnosis and incorporate relevant clinical terms into prompts for GPT-3.5 and -4.
A notable trend is the model selection between open-source and closed-source LLMs, such as Llama 4 vs GPT-4. Approximately half of the reviewed studies (67.8%, $n=40$) directly utilize powerful closed-source models such as GPT-3.5 and GPT-4 through proprietary APIs. This approach prioritizes rapid development and sophisticated output quality without extensive model customization. In contrast, 39.0% ($n=23$) of prompt-based studies utilize fine-tuned open-source models (e.g., GPT-2 [68, 19, 25, 22, 24, 20], T5 [37, 8]), offering flexibility to specialize the model with biomedical domain-specific data. Fine-tuning [26, 17, 24] involves additional training of a pretrained model on a smaller, task-specific dataset, enabling the model to specialize its outputs according to specific biomedical contexts or data types. Fine-tuning open-source models demonstrates advantages in privacy-sensitive or -nuanced clinical contexts [20, 26, 22], such as assisted suicide detection [8] and cancer diagnosis [45]. For example, Lu et al. [22] fine-tuned GPT-2 on MIMIC-III [70] discharge summaries to generate synthetic clinical notes for data augmentation, effectively mitigating class imbalance, preserving patient privacy, and improving the accuracy of a downstream readmission-prediction model. However, while open-source models offer greater customization potential, they typically underperform closed-source models in general language tasks and clinical reasoning benchmarks [57, 34, 62, 8, 12, 36]. For example, Jeong et al. [36] report that the open-source model LLaMA2 7B underperforms GPT-4-base on clinical reasoning benchmarks.
[…]

### v2

@@ 2 Results
Table 1: Summary of data types, generation methods, accessibility, purposes and medical applications in the collected studies. Languages include English (EN), Dutch (NL), French (FR), Chinese (ZH), and Arabic (AR). Generation methods comprise fine-tuning (SFT: supervised, DAFT: domain-adaptive, TAFT: task-adaptive, IT: instruction tuning, None) and prompting (ZS: zero-shot, FS: few-shot, INST: instruction-driven, CoT: reasoning-augmented, RAG: knowledge-augmented). Synthetic data purpose includes Training (used independently), Supplement (used with real data for joint training), and Privacy (used for privacy preservation such as sharing or de-identification). Data accessibility is classified as Yes (publicly available), Request (available upon request), and No (not publicly available or not specified). Medical applications cover question answering (QA), information extraction (IE), and social determinants of health (SDoH).
Study Modality Lang. Year Generative Model Approach Accessible Purpose Medical Application Fine-tuning Prompting [shakeri-etal-2020-end] Text EN 2020 BART TAFT INST No Training Biomedical QA [ive2020generation] Text EN 2020 Transformer SFT INST No Training Phenotype inference [li2021synthetic] Text EN 2021 GPT-2 TAFT ZS Yes Training Clinical IE [libbi2021generating] Text NL 2021 GPT-2 DAFT ZS No Training De-identification [chintagunta-etal-2021-medically] Text EN 2021 GPT-3 None FS, RAG No Training Dialogue summarization [lu2022textual] Text EN 2022 GPT-2 TAFT ZS No Training Readmission prediction [peng2023study] Text EN 2023 GPT-3 None ZS Yes Training General [hiebel-etal-2023-synthetic] Text FR 2023 GPT-2, Bloom DAFT ZS Yes Training Clinical IE [khademi2023data] Text EN 2023 GPT-2 TAFT ZS No Training Syndromic detection [spitale2023exploring] Text EN 2023 GPT-3.5 IT INST Yes Training Assisted suicide [theodorou2023synthetic] Tabular EN 2023 HALO TAFT INST Yes Training General [li2023two] Text EN 2023 Llama, GPT-4 None CoT Request* Training Alzheimer’s disease [sufi2024addressing] Text EN 2024 ChatGPT API None INST Yes Training Discharge summary [wu2024callm] Text EN 2024 GPT-4 None INST, CoT Request* Training PTSD diagnosis [chen2024chatffa] Image ZH, EN 2024 GPT-3.5 None INST No Training Medical imaging [litake2024constructing] Text EN 2024 GPT-3.5, Llama-2 None INST Yes Training Acute renal failure [sarkar2024identification] Text EN 2024 GPT-3.5 None INST No Privacy Phenotype inference [nievas2024distilling] Text EN 2024 GPT-4 None INST, CoT No Training Cohort selection [wang2024dke] Text EN 2024 GPT-3.5 None INST, RAG No Supplement Cohort selection [moser2024generating] Text DE 2024 Zephyr, GPT-4 None INST, RAG No Training Emergency medicine [bird2024generative] Text EN 2024 Mistral None ZS, INST Yes Training Psychological therapy [jeong2024improving] Text EN 2024 GPT-4, Llama-2 IT ZS, FS Yes Training Medical QA [zafar2024ki-mag] Text EN 2024 T5 TAFT RAG Request Training Medical QA [xu-etal-2024-knowledge] Text EN 2024 GPT-3.5 None FS, RAG Yes Training General [guevara2024large] Text EN 2024 GPT-3.5/4 None FS, INST Yes Supplement SDoH extraction [ehrett2024leveraging] Tabular EN 2024 ChatGPT None FS No Supplement Medical education [wang2024notechat] Text EN 2024 GPT-3.5 None INST, CoT Yes Supplement Dialogue generation [gabriel2024development] Text EN 2024 GPT-3.5 None ZS, INST No Training SDoH extraction [gao2024pressinpose] Tabular EN 2024 GPT-4 None FS No Training Gait rehabilitation [kweon2024publicly] Text EN 2024 GPT-3.5 None FS, INST Yes Training General [weerasinghe2024real] Audio EN 2024 GPT-4 None INST Yes Supplement EMS assistance [delmas2024relation] Text EN 2024 Vicuna None INST Yes Training Clinical IE [ghanadian2024socially] Text EN 2024 GPT-3.5, Flan-T5, Llama-2 None ZS, FS No Supplement Suicide prediction [zeinali2024symptom] Text EN 2024 GPT-4 None INST No Training Cancer symptom extraction [mishra2024synthetic] Text EN 2024 ChatGPT None INST Yes Training Clinical summary [zhang-etal-2024-synthetic] Text EN 2024 GPT-3.5 None INST, RAG Yes Training Medical QA [dobhal2024synthetic] Tabular - 2024 GPT-4o None FS, CoT No Supplement Nursing care [wang2024taxonomy] Text EN 2024 GPT-4 None INST, ZS Yes Training Patient portal triage [jones2024teaching] Text EN 2024 Vicuna v1.1, Orca IT INST No Training Report generation [alghamdi2024towards] Text AR 2024 ChatGPT None INST, RAG Yes Supplement Health chatbot [wang2024twin] Tabular EN 2024 ChatGPT None FS, INST No Supplement Clinical trial simulation [bucur2024leveraging] Text EN 2024 Llama, Alpaca None ZS, INST No Supplement Depression symptom [woolsey2024utilizing] Text EN 2024 GPT-3.5/4 None ZS, INST No Supplement Autism diagnosis [chen2025automated] Text EN 2025 ChatGPT, Llama, Gemini None INST Request Privacy Stroke prediction [scroggins2025does] Audio EN 2025 GPT-4 None FS, CoT No Training Dialogue analysis [albayrak2025enhancing] Text EN 2025 GPT-3.5 None ZS, INST No Supplement Phenotype ontology [liu2025generating] Text EN 2025 GPT-3.5/4, Llama-2, Mistral None ZS, FS No Training Radiology prediction [wang2025hypnos] Text ZH 2025 GPT-3.5, Claude None FS, RAG Request Training Anesthesia QA [theodorou2025improving] Tabular - 2025 HALO SFT INST Yes Supplement Patient outcome prediction [cai2025improving] Text ZH 2025 ChatGPT None FS, INST Yes Training Mental health IE [barr2025large] Tabular EN 2025 GPT-4o None ZS, INST Yes Privacy Data scarcity [chuang2025robust] Text EN 2025 GPT-3.5/4, Mistral None FS, INST No Privacy Phenotype inference [kim2025small] Text EN 2025 GPT-4 None CoT Request* Training Medical reasoning [li2025synthetic] Text ZH 2025 GPT-3.5 None INST Request Training Database search [ashofteh2025targeted] Text EN 2025 Llama3-instruct None FS No Training Metastases detection [ejiga2025text] Image EN 2025 Stable Diffusion None FS Yes Supplement Colonoscopy analysis [li2025topofm] Image EN 2025 ChatGPT API None INST No Supplement Pathology analysis [vsuvalov2025using] Text ET 2025 GPT-2 TAFT ZS, FS Request Training Clinical IE [miletic2025utility] Tabular EN 2025 DistilGPT-2 SFT INST No Training Patient outcome prediction Notes. A dash (–) indicates unknown information. The “Request*” indicates data were stated to be publicly available but are either not yet released or temporarily inaccessible. Model names are reported as in the original papers; when only “ChatGPT” or “ChatGPT API” was mentioned, we list it accordingly.
2.1 Synthetic Data Types and Medical Applications
Increasing model developments for biomedical applications lead to the surging needs of synthetic data [ive2020generation, libbi2021generating]. In this section, we provide an overview of recent studies using synthetic data and address three critical questions: 1) What types of synthetic data are being generated? 2) What modalities and clinical tasks do these datasets target? 3) Are those data resources accessible for reusable research, such as benchmarking and model training? To answer those questions, we examine several aspects of the collected studies and summarize them in Table 1, including data modality, data language, data size, published year, medical application, and data or model accessibility. We aim to provide insights into current and future trends of data types and biomedical applications.
Biomedical synthetic data has shown diverse medical topics, languages, and modalities. Most of the studies (55.9%, $n=33$) are for general topics, while others include specific medical areas, such as suicide [spitale2023exploring, ghanadian2024socially], colorectal cancer [ejiga2025text], and radiology [liu2025generating]. For example, studies generate synthetic data for augmenting diagnosis accuracy for Post-traumatic stress disorder (PTSD) [wu2024callm] and Autism Spectrum Disorders (ASD) [woolsey2024utilizing]. English synthetic data accounted for the majority of generated datasets (84.7%, $n=50$). Synthetic data were also generated in French [hiebel-etal-2023-synthetic], Chinese [chen2024chatffa, wang2025hypnos, cai2025improving, li2025synthetic], German [moser2024generating], and Arabic [alghamdi2024towards], demonstrating initial efforts to support non-English language resources. For example, Moser et al. [moser2024generating] employed multilingual LLMs to generate German emergency-medical dialogues simulating realistic interactions between ambulance staff and patients. Most datasets focus on unstructured text modality (e.g., clinical narratives [wang2024notechat, chintagunta-etal-2021-medically, jones2024teaching] and discharge summaries [sufi2024addressing]), while a smaller but growing subset includes tabular [theodorou2023synthetic, ehrett2024leveraging, gao2024pressinpose, dobhal2024synthetic, wang2024twin, theodorou2025improving, barr2025large, miletic2025utility], image [chen2024chatffa, ejiga2025text, li2025topofm], and audio [scroggins2025does]. For example, Theodorou et al. [theodorou2023synthetic] generate novel tabular datasets of high-dimensional longitudinal EHR records to provide realistic, privacy-preserving alternatives for machine learning and statistical analysis, while Ejiga et al. [ejiga2025text] generate synthetic colonoscopy image datasets via fine-tuned text-to-image synthesis to augment training data for robust colorectal cancer detection and precise polyp segmentation. Data sizes vary across health issues (e.g., cancer and mental health), yet the majority has relatively small sets with fewer than 10K samples, indicating the critical utility of synthetic data under low-resourced scenarios. For example, FairPlay [theodorou2025improving] generates synthetic data from authentic EHR records to bolster under-represented patient subgroups, boosting mortality‐prediction F1 scores by up to 21% and markedly narrowing performance gaps across different demographic groups. The various synthetic datasets highlight a growing community to deploy the biomedical synthetic data.
Synthetic data generation in biomedical research increasingly supports a wide array of clinical applications. Those tasks cover diverse medical needs, including phenotype classification [zeinali2024symptom, woolsey2024utilizing, khademi2023data, li2023two, ashofteh2025targeted], PTSD symptom extraction [wu2024callm], de-identification (SynNote) [libbi2021generating], clinical note summarization [sufi2024addressing, mishra2024synthetic], and mental health assessment [bucur2024leveraging, bird2024generative, ghanadian2024socially]. For example, CALLM [wu2024callm] constructs clinical interview datasets to aid PTSD diagnosis by generating synthetic transcripts; and Barabadi et al. [ashofteh2025targeted] generated synthetic radiology reports to enable automatic detection of metastatic sites in cancer patients. Those applications broaden the utility of synthetic data for health challenges, such as data shortage [wu2024callm] and privacy concerns [libbi2021generating, sarkar2024identification, chuang2025robust, barr2025large, chen2025automated]. However, accessibility remains a central concern for advancing synthetic data utility (see Supplementary Table S1 for a detailed list of experimental datasets and their corresponding links). Table 1 shows 50.8% ($n=30$) of the studies are explicitly described as accessible, providing open or partially open resources for reproducible development, while others’ accessibility details or licensing remain unclear. Open datasets such as Syn-HPI [li2021synthetic], ClinGen [xu-etal-2024-knowledge], and Syn-EMS-Audio [weerasinghe2024real] serve as synthetic data benchmarks for evaluating new models and methods, while others—particularly those derived from sensitive clinical domains—may be restricted or anonymized to protect patient privacy. Ensuring clear, well-documented access protocols and licenses will be essential for fostering broader collaboration and translational impacts of synthetic data in biomedical research.
2.2 Synthetic Data Generation Methodology
Emerging Transformer-based large language models (LLMs) have significantly advanced synthetic data generation methodologies in biomedical research, enabling diverse applications, as summarized in Table 1. These novel approaches address challenges associated with limited clinical data by augmenting datasets, thus enhancing generalizability and mitigating overfitting in downstream machine learning tasks across various healthcare domains. In this section, we systematically examine generation methodologies among 59 reviewed studies, summarizing recent trends in model architectures, generation techniques, and integration strategies for synthetic data generation. Specifically, we aim to answer two key questions: 1) what predominant methods (e.g., prompt-based vs. specialized architectures) are currently employed for biomedical synthetic data generation; and 2) what trends and variations do exist regarding model selection (open-source vs. closed-source) and prompting strategies (zero-shot, few-shot, reasoning-augmented)? Because several studies employed combinations of approaches (e.g., both open- and closed-source models, or few-shot with CoT or RAG), each study was counted in all applicable categories. In the case of prompting-based methods, studies employing multiple prompting strategies were summarized based on their two predominant prompting types for alignment, and the overlaps among these categories are visualized in Figure 2.
Figure 2: Overlap among prompting strategies (INST, Zero-shot, Few-shot, CoT, RAG) in the reviewed studies. Each intersection represents studies combining multiple methods.
Our review identifies prompt-based generation as the predominant synthetic data generation method, employed by approximately 74.6% ($n=44$) of the analyzed studies. Prompt-based approaches primarily rely on meticulously crafted textual instructions, leveraging fine-tuned LLMs, such as T5 and GPT variants (e.g., GPT-4) for diverse biomedical tasks [li2025synthetic, litake2024constructing, wang2024twin]. The most prevalent strategy is instruction-driven (INST) prompting, observed in 42.4% ($n=25$) of studies, which guides generation through detailed, explicit instructions that typically encompass core task definitions, persona assignments, and specific constraints on the desired content, format, and style. It leverages the model’s ability to follow complex commands without in-context examples, making it highly versatile. For instance, Sufi et al.[sufi2024addressing] utilized detailed instructions to generate realistic discharge summaries, while Wang et al.[wang2024taxonomy] employed structured prompts to synthesize patient portal messages for creating a corpus to analyze linguistic features and support downstream applications like patient portal message triage. Zero-shot prompting represents another approach, used in 28.8% ($n=17$) of studies, involving prompting LLMs without illustrative examples and relying solely on explicit task instructions and the pretrained knowledge encoded in LLMs [barr2025large]. This method is suitable for simpler tasks, such as dialogue summarization [chintagunta-etal-2021-medically] or generating structured medical reports [jones2024teaching], where extensive contextual training is less critical. For example, Barr et al. [barr2025large] used zero-shot prompting with GPT-4o to synthesize perioperative clinical tables of patient demographics, surgical parameters and outcomes, demonstrating that the model can produce realistic synthetic data and preserve key statistical patterns based solely on qualitative instructions. Few-shot prompting provides a small number of curated examples (typically 2-5) within the prompt to illustrate the task and guide LLMs for synthetic data generation, which accounts for 20.3% ($n=12$) of studies. This approach uses limited curated examples to guide LLMs in generating clinically relevant synthetic data across diverse applications, including suicide symptom extractions [ghanadian2024socially], phenotype classification [zeinali2024symptom, woolsey2024utilizing, khademi2023data, li2023two, ashofteh2025targeted], and therapeutic dialogue generation [wang2024notechat, bird2024generative]. For example, Ghanadian et al.[ghanadian2024socially] use a patient’s suicidal narrative and a medication counseling dialogue as few-shot prompts to generate clinical texts for suicide symptom extraction and therapeutic dialogue generation, while Nievas et al.[nievas2024distilling] use clinical trial inclusion and exclusion criteria paired with patient summaries as few-shot prompts to generate a synthetic patient–trial matching dataset. Beyond providing instructions and examples, more sophisticated techniques aim to enhance the reasoning and factuality of generated data. Reasoning-augmented prompting, commonly known as Chain-of-Thought (CoT) and used in 10.2% ($n=6$) of studies, enhances generation quality by instructing the model to first articulate its step-by-step reasoning process before producing the final output. This method improves the coherence and logical validity of synthetic data for complex clinical tasks. For example, Wu et al. [wu2024callm] applied CoT to generate synthetic PTSD diagnostic interviews by prompting the model to provide both a simulated patient’s response and the clinical reasoning for it, ensuring the dialogue aligned with DSM-V criteria, while Kim et al.[kim2025small] used CoT to improve medical reasoning for question-answering tasks. Furthermore, retrieval-augmented generation (RAG), improves factual accuracy by dynamically incorporating information from external knowledge bases into the prompt. This technique, found in 11.9% ($n=7$) of studies, retrieves relevant documents (e.g., from clinical guidelines or medical literature) to ground the generation in verifiable, up-to-date information. For instance, Wang et al.[wang2024dke] employed a RAG-like method for data augmentation, retrieving semantically similar synonyms from a knowledge graph to replace key terms. Similarly, Zafar et al.[zafar2024ki-mag] enhanced medical QA by grounding generated answers with knowledge dynamically retrieved from the UMLS.
A notable trend is the model selection between open-source and closed-source LLMs, such as Llama vs GPT-4. Approximately 67.8% ($n=40$) of the reviewed studies directly utilize closed-source models such as GPT-3.5 and GPT-4 through proprietary APIs. This approach prioritizes rapid development and output quality without extensive model customization. The remaining 32.2% ($n=19$) of studies employ fine-tuning approaches on open-source models, which involves additional training of pretrained models on domain-specific datasets to adapt them for particular biomedical contexts. Fine-tuning approaches can be categorized into several distinct types based on their adaptation strategies. Supervised fine-tuning (SFT) involves training models on labeled, task-specific datasets where input-output pairs are explicitly provided [ive2020generation, theodorou2025improving, miletic2025utility]. Ive et al. [ive2020generation] demonstrated this approach by fine-tuning Transformer models on real EHRs using a gap-filling task, where models learned to generate complete clinical documents conditioned on key phrases and clinical metadata for phenotype classification tasks. Domain-adaptive fine-tuning (DAFT) adapts models to broad clinical domains by training on large corpora of domain-specific text without task-specific labels [libbi2021generating, hiebel-etal-2023-synthetic]. This approach helps models acquire clinical vocabulary, medical terminology, and domain-specific linguistic patterns. Libbi et al. [libbi2021generating] employed DAFT by fine-tuning an English GPT-2 model on Dutch clinical narratives with explicit PHI annotations, enabling joint generation of synthetic text and de-identification labels for training downstream NER models. Task-adaptive fine-tuning (TAFT) focuses on specific downstream tasks by training models on task-relevant data [shakeri-etal-2020-end, li2021synthetic, lu2022textual, khademi2023data, theodorou2023synthetic, zafar2024ki-mag, vsuvalov2025using]. Lu et al. [lu2022textual] applied TAFT to GPT-2 by prepending class labels to MIMIC-III discharge summaries during training, enabling the model to generate synthetic clinical notes with embedded readmission status for downstream prediction tasks. Instruction tuning (IT) trains models to follow structured commands and detailed instructions, enhancing their ability to generate content according to specific requirements [spitale2023exploring, jeong2024improving, delmas2024relation, jones2024teaching]. For instance, Jeong et al. [jeong2024improving] employed IT to fine-tune LLaMA2 models on medical instruction datasets, developing critic models that annotate and filter synthetic medical instructions to generate instruction sets for training generator models. Fine-tuned open-source models demonstrate particular advantages in privacy-sensitive clinical contexts and specialized medical domains. However, these models typically require more computational resources and domain expertise compared to direct API usage of closed-source models. Studies comparing model performance indicate that while fine-tuned open-source models offer greater customization potential, they generally underperform closed-source models on standardized clinical reasoning benchmarks. For example, Jeong et al. [jeong2024improving] observed that fine-tuned LLaMA2 7B models achieved lower performance than GPT-4 on clinical reasoning tasks, despite domain-specific adaptation.
Approximately 25.4% ($n=15$) of the reviewed studies fine-tuned pre-trained LLMs or explore specialized model architectures for synthetic data generation, including multimodal, multi-agent frameworks and custom-designed transformer architectures (e.g., HALO) [li2025topofm, ejiga2025text, weerasinghe2024real, scroggins2025does, gao2024pressinpose, wang2024notechat, theodorou2025improving, theodorou2023synthetic, ive2020generation]. Multimodal approaches integrate multiple data modalities, such as text, images, and audio, to create richer, more realistic synthetic datasets, such as colonoscopy image analysis [ejiga2025text], digital pathology interpretation [li2025topofm], and emergency medical services (EMS) assistance [weerasinghe2024real]. For example, CognitiveEMS [weerasinghe2024real] prompts an LLM with emergency protocols to generate synthetic cardiac-arrest dialogues, converts them to speech, and synchronizes the audio with AR smart-glasses video frames annotated by a zero-shot vision classifier, producing a unified dataset for cardiac emergency analysis. Multi-agent frameworks [alghamdi2024towards, wu2024callm, wang2024notechat] involve employing multiple interacting models or agents, each performing distinct roles, to collaboratively produce complex and realistic data, such as doctor-patient dialogues or clinical scenarios. For instance, NoteChat [wang2024notechat] employs a cooperative multi-agent framework with planning, role play, and refinement modules to generate synthetic patient–physician dialogues that mirror real clinical encounters. Custom transformer architectures[theodorou2025improving, theodorou2023synthetic, ive2020generation] extend the standard decoder by integrating domain-specific metadata and employing hierarchical modeling strategies, first capturing broader structural information and then refining finer details, to generate structured and realistic synthetic clinical datasets tailored to specific medical scenarios. For example, Theodorou et al. [theodorou2023synthetic] demonstrate that their hierarchical autoregressive model generates privacy-preserving EHRs whose temporal and code distributions yield downstream prediction performance on par with real data. Collectively, these specialized frameworks enhance the fidelity and applicability of synthetic biomedical data for various complex health scenarios.
Figure 3: Evaluation characteristics of reviewed studies: disease, downstream task, data size, number of automated metrics (Metrics #), inclusion of intrinsic evaluation (Intrinsic+), human involvement in generation or task evaluations (Human-in-the-Loop), and LLM-based evaluation (LLM Eval). CLS, NER, IE, RE, QA, NLI, and ASR refer to downstream tasks of classification, named entity recognition, information extraction, relation extraction, question answering, natural language inference, and automatic speech recognition, respectively.
[…]